OpenStudy (luckycoins888):

solve for y: 5x - 15y = 20

OpenStudy (luckycoins888):

@poopsiedoodle

OpenStudy (elonasushchik):

5x+15y=20 x+3y=3 is not solvable 5x+ 15y = 20 -5(x + 3y)= -5(3) 5x+ 15y = 20 -5x - 15y= -15 0 = 5 so it's false for all x and y Hope this helps..